# DyMO: Training-Free Diffusion Model Alignment with Dynamic Multi-Objective Scheduling

### v2

@@ 5.2.2 Qualitative Evaluation
We compare our method against a baseline model (SD V1.5 [31]), training-free models (DNO [40], FreeDom [53]) and training-based models (AlignProp [29], Diffusion-DPO [44], Diffusion-KTO [22], SPO [23]), as illustrated in Fig. 3. It is observed that our method contains coherent layout and rich semantic information aligned with the prompt, while achieving overall visual appeal and aesthetic quality. Additionally, we apply our framework to SDXL [28], Diffusion-DPO [44], SPO [23] and compare some sample image generation results in Fig. 4. The results show that there is a significant improvement in image quality before and after using our method. Specifically, our method enhances the color vibrancy and detail richness, bringing out more defined textures on fur and fabric as well as vivid lighting effects. More visual comparison results are provided in the supplementary material.
Figure 4: Qualitative comparison based on SDXL backbones. The text prompts used to generate the images are provided in the Tab. 7.
@@ 5.2.3 Case Performance
(b) Human evaluation with and without our method on PartiPrompts and HPSv2 datasets.
@@ 5.3 Ablation Study
Effect of layout loss $\mathcal{L}_{A}$. The structural formation of image content occurs early in the denoising process, as verified by Tab. 3. We hypothesize that the inferior performance is due to the loss of semantic components in the prompt.
@@
Effect of preference loss $\mathcal{L}_{R}$. The results in Tab. 3 highlight the importance of the human feedback. Accurate preference guidance enhances the image-prompt alignment and improves overall image quality. Additionally, we use the PickScore [19] to replace the step-aware preference model [23], achieving improvements over the baseline model, which further demonstrates the effectiveness and versatility of our method.
@@
Effect of dynamic recurrent strategy. We further conduct experiments with fixed iteration counts set to 1, 5, and 10 to verify the effectiveness of our method, as presented in Tab. 4. Our approach achieves the better performance with a moderate time cost, demonstrating that the dynamic recurrent strategy effectively meets the need of diffusion models.
@@ 8.5 Additional Results on Dynamic Time Travel Steps
In Tab. 4 in the main paper, we provide ablation study results of different time-travel steps. By comparison, our proposed dynamic recurrent strategy effectively achieves better performance on different metrics. We provide some visual comparison examples in Fig. 7, where we can observe that the visual qualities of the generated image are consistent with the numerical results. The proposed method with dynamic recurrent step scheduling is effective in producing better results with less time. With the proposed alignment objectives, the proposed method can also work well with less (and fixed) recurrent step numbers. In addition, compared with the other training-free methods with full-chain backpropagation methods, like DNO [40] and PromptOpt [7], require 370 and 280 seconds, respectively, our method can be more efficient and effective. Compared with the one-step approximation methods, such as FreeDom [53] (with 170 seconds), our method is also efficient and performs better on the results. While many training-free methods take more time to guide the denoising process, our method can also achieve efficient image quality improvement in less time by employing fewer time-travel steps, as demonstrated in Tab. 4 and Fig. 7. The whole generation process can be further accelerated by incorporating more efficient sampling processes, which is left as future work.

### v3

@@ 5.2.2 Qualitative Evaluation
We compare our method against a baseline model (SD V1.5 [32]), training-free models (DNO [41], FreeDom [54]) and training-based models (AlignProp [30], Diffusion-DPO [45], Diffusion-KTO [23], SPO [24]), as illustrated in Fig. 3. It is observed that our method contains coherent layout and rich semantic information aligned with the prompt, while achieving overall visual appeal and aesthetic quality. Additionally, we apply our framework to SDXL [29], Diffusion-DPO [45], SPO [24] and compare some sample image generation results in Fig. 4. The results show significant improvement in image quality before and after using our method. Specifically, our method enhances the color vibrancy and detail richness, bringing out more defined textures on fur and fabric as well as vivid lighting effects.
Figure 4: Qualitative comparison based on SDXL backbones.
@@ 5.2.3 Case Performance
(b) Human evaluation w/ & w/o our method on PartiPrompts & HPSv2.
@@ 5.3 Ablation Study
Effect of semantic alignment loss $\mathcal{L}_{A}$. The structural formation of image content occurs early in the denoising process, as verified by Tab. 3. The inferior performances are due to the lost of semantic components in the prompt.
Effect of preference alignment loss $\mathcal{L}_{R}$. The results in Tab. 3 highlight the importance of the human feedback. Accurate preference guidance enhances the image-prompt alignment and improves overall image quality. Additionally, we use the PickScore [20] to replace the step-aware preference model [24], achieving improvements over the baseline model, which further demonstrates the effectiveness and versatility of our method.
@@
Effect of open-source LLMs. DyMO’s LLM-based text prompt pre-analysis is moderate for LLMs and does not require advanced models (LLM-agnostic). We validate it with open-source Llama3.30-70B (Tab. 3), achieving strong performance (22.58)—better than others, slightly below GPT-4-based results (23.07).
Effect of dynamic recurrent strategy. We conduct experiments with fixed iteration counts set to 1, 5, and 10 as presented in Tab. 4. Our approach achieves the better performance with a moderate time cost, demonstrating the effectiveness of dynamic recurrent strategy.
@@ 8.3 Additional Qualitative Results
Table 5: Geneval Benchmark evaluation based on SD V1.5.
Methods Overall Single object Two object Counting Colors Position Color attribution
SD V1.5 0.42 1.00 0.38 0.35 0.77 0.04 0.00
DNO 0.43 0.96 0.35 0.35 0.82 0.04 0.05
PromptOpt 0.39 0.96 0.25 0.23 0.80 0.04 0.05
FreeDom 0.52 1.00 0.56 0.64 0.80 0.02 0.13
AlignProp 0.42 1.00 0.23 0.41 0.74 0.04 0.06
Diffusion-DPO 0.47 1.00 0.48 0.47 0.80 0.02 0.05
Diffusion-KTO 0.49 1.00 0.53 0.41 0.82 0.06 0.11
SPO 0.47 0.96 0.48 0.47 0.74 0.11 0.05
SD V1.5+Ours 0.57 1.00 0.72 0.47 0.83 0.07 0.34
@@
8.4 Additional Quantitative Results
We conduct quantitative evaluation on Geneval benchmark [14] and show comparisons in Tab. 5. Our method performs very well and shows superiority in many aspects, e.g., overall, attribute binding and object synthesis. The proposed semantic alignment can help DyMO on multi-object synthesis. Tab. 5 also shows that DyMO performs well in multiple objects and counting.
@@ 8.5 Additional Details on Human Evaluation
(a) PickScore
(b) HPSv2
(c) ImageReward
(d) Aesthetics
Figure 9: Insensitivity to scheduling parameter $t_{1}$, $t_{2}$.
8.6 Additional Results on Dynamic Scheduling
The scheduling weights (Eq. 10) are automatically adjusted based on content changes to balance the roles of $\mathcal{L}_{A}$ and $\mathcal{L}_{R}$ in different stages. In the early stage, $\mathcal{L}_{A}$ dominates due to content instability and noise (where $\mathcal{L}_{R}$’s guidance is weak). As semantics stabilize, $\mathcal{L}_{R}$ gains more weights, refining alignment with more detailed visual preferences. Thus, the method prioritizes semantic alignment first, then preference alignment. Based on the observed dynamics of $w_{R}$ and $w_{A}$ (trend and smoothness), we set the stage split $(t1,t2)$ as $(800,500)$ for efficiency and simplicity. The model is insensitive to this hyperparameter, as confirmed by the grid-search analysis (Fig. 9).
@@ 8.7 Additional Results on Dynamic Time Travel Steps
In Tab. 4 in the main paper, we provide ablation study results of different time-travel steps. By comparison, our proposed dynamic recurrent strategy effectively achieves better performance on different metrics. We provide some visual comparison examples in Fig. 7, where we can observe that the visual qualities of the generated image are consistent with the numerical results. The proposed method with dynamic recurrent step scheduling is effective in producing better results with less time. With the proposed alignment objectives, the proposed method can also work well with less (and fixed) recurrent step numbers. In addition, compared with the other training-free methods with full-chain backpropagation methods, like DNO [41] and PromptOpt [7], require 370 and 280 seconds, respectively, our method can be more efficient and effective. Compared with the one-step approximation methods, such as FreeDom [54] (with 170 seconds), our method is also efficient and performs better on the results. While many training-free methods take more time to guide the denoising process, our method maintains strong performance in just 40s with fewer iterations as demonstrated in Tab. 4 and Fig. 7. The whole generation process can be further accelerated by incorporating more efficient sampling processes, which is left as future work.